Students can buy tickets to a basketball game for $1. The admission for nonstudents is $2. If 350 tickets are sold and the total receipts are $450, how many student tickets are sold?
Answer by mananth(16949)

Students can buy tickets to a basketball game for $1. The admission for nonstudents is $2. If 350 tickets are sold and the total receipts are $450, how many student tickets are sold?
...
let student tickets sold be x--------------$1
non student will be 350-x---------------$2
..
1*x+2(350-x)=450
x+700-2x=450
-x=-250
x= 250 student tickets
..
m.ananth@hotmail.ca